A certain weak acid, HA, has a Ka value of 2.0×10−7.

Part A Calculate the percent dissociation of HA in a 0.10 M solution

Part B Calculate the percent dissociation of HA in a 0.010 M solution
Answer #1

Equation:
Ka = [H+]² /[HA]
2.0*10^-7 = [H+]² / 0.10
[H+]² = (2.0*10^*-7)*0.1
[H+]² = 2.0*10^-8
[H+] = √(2.0*10^-8)
[H+] = 1.414*10^-4

% dissociation = (1.414*10^-4) / 0.10 *100 = 0.1414% dissociated.

Part B ; For 0.01M solution
Ka = [H+]² / [HA]
2.0*10^-7 = [H+]²/0.01
[H+]² = (2.0*10^-7)*0.01
[H+]² = 20.0*10^-10
[H+] = 4.4721*10^-5
% dissociation = (4.4721*10^-5) /0.01*100 = 0.44721% dissociated.